OpenStudy (anonymous):

What's the equation of the line passing through (-1,-9) and (3,11) , in the form y=mx+c?

OpenStudy (anonymous):

that points lie on the line .. plug them in the equation -9=-m+c 11=3m+c solve for m and c

OpenStudy (anonymous):

Can also find m first using slope formula \[m=\frac{y_2-y_1}{x_2-x_1} \space .\] Then use one point and y=mx+c to solve for c.